My mother is 35 years old
She is that age without counting the weekends each year
What is her real age!
Please Help !!
Thank You
xxx

Answer by stanbon(75887)

My mother is 35 years old
She is that age without counting the weekends each year
What is her real age!
Since the weekends comprise 2/7 ths of the week, in
a year they comprise 2/7 of a year.
EQUATION:
Your mother is 35 +(2/7)35= 35+10= 45 yrs old.
Cheers,
Stan H.